OpenStudy (anonymous):

The formula for degrees Celsius is: C =5/9(F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 20 degrees Celsius is.

OpenStudy (anonymous):

5=C*(9(F-32)) 5/C=(9(F-32)) 5/C*9=(F-32) F=(5/C*9)+32

OpenStudy (anonymous):

5/9(20-32) 5/9*12 5/108

OpenStudy (anonymous):

im still confused. can you decribe the steps as you go please?

OpenStudy (anonymous):

\[F=(9/5)C+32=1.8C+32\] 1.8 x (20 Celsius)+32=68 Farenheit

OpenStudy (klimenkov):

Look, you want to find \(F\) from the equation \(C=\frac59(F-32)\). 1 step: Multiplying both parts of the equation on \(\frac95\): \[\frac95C=\frac95\cdot\frac59(F-32)\]We have that \(\frac95\cdot\frac59=1\). So, we have \[\frac95C=1\cdot(F-32)\]\[\frac95C=F-32\]2 step: Adding 32 to both parts of the equation: \[\frac95C+32=F-32+32\]We have that \(-32+32=0\). So, we have \[\frac95C+32=F+0\]\[\frac95C+32=F\]Or \(F=\frac95C+32\).

OpenStudy (klimenkov):

Do you understand this?